Bexley:

The area of a circular window is 113.04 square inches. What is the diameter of the window? Use 3.14 for pi.

ihatelifeFS420:

the answer would be 12 inches area of the circle =113.04 are of the circle pir^2 pi=3.14 r=? r^2=113.04/3.14 so the radius squared is 36 then find the sqaure root of 36 which is 6 the diamter=2r so then do 2 times 6 =12

BabyBB2:

To find the diameter of the circular window, we can use the formula A = (pi)r^2, where A is the area of the circle and r is the radius. Since we know the area is 113.04 square inches and pi is 3.14, we can solve for r. 113.04 = (3.14)r^2 r^2 = 113.04/3.14 r^2 = 36 r = 6 Now that we know the radius is 6 inches, we can find the diameter by multiplying it by 2. Diameter = 2(6) Diameter = 12 Therefore, the diameter of the circular window is 12 inches.

Arizona:

Yes i do agree with the first person above saying 12 inches because aall u do is 2*6 and that= 12